Giving the following data for XY Company

year

2008

2009

Total Sales

$145,0000

$121,4000

Cost of goods sold

$880,000

$738,400

Selling & Administrative Expenses

$92,500

$80,700

Assume high-low method is adequate in this case, the total fixed costs of selling and administrative expenses is

1.

$10,000

2.

$12,000

3.

$15,000

4.

$20,000
Answer #1

Variable rate for selling and administrative expense = (92500-80700)/(1450000-1214000) = 0.05

Fixed cost = Total Cost-Variable cost = 92500-(1450000*.05) = 20000

So answer is 4) $20000